In an isosceles triangle, the ratio of the measure of each base angle to the vertex angle is 2:5. Find the measure of each angle of the triangle.
Answer by ikleyn(52787)

.

The vertex angle is 2x;


each of the two base angles is 5x.



The sum of the angles equation is

    2x + 5x + 5x = 180  degrees,   or


        12x      = 180,

          x      = 180/12 = 15.



ANSWER.  The angles are  2*15 = 30 angles (the vertex);

         5*15 = 75 degrees  (each of the two base angles).

Solved.